Naciama:

Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. that has an extraneous solution

Naciama:

Please I need help I only have limited time

SmokeyBrown:

The question mentions an equation that you should use as a model. May we see this equation? It seems like an important part of the question.

Naciama:

They don't give a model which is weird but you can make your own

SmokeyBrown:

Ah. Interesting. In general, an "extraneous solution" exists when an equation can be manipulated into a new form, which produces a solution that does not satisfy the original. That's a bit abstract, but there is a more concrete example here that should help you understand: https://www.onlinemath4all.com/extraneous-solutions-of-radical-equations.html#:~:text=Extraneous%20solutions%20of%20radical%20equations%20%3A%20Squaring%20each,the%20original%20equation%20when%20you%20solve%20radical%20equations. Does the question perhaps give you a template that you're supposed to make your equations off of?

Naciama:

√(x-2) = (x- 4) something like this

Naciama:

Im in algebra 1

SmokeyBrown:

Interesting. I'm not sure if it's a coincidence, but the article I linked actually gives an example of how √(x-2) = (x- 4) can produce an extraneous solution. Basically, √(x-2) = (x- 4) can be converted to x - 2 = x^2 - 8 x + 16 by squaring both sides of the original. This new equation can be satisfied by 6 and 3, but only 6 satisfies the original, which makes 3 an extraneous solution. So if you have that, it seems like you're halfway done! Now we just need to write an equation that does *not* have any extraneous solutions.

Naciama:

I was wondering if I could just get the answers if your okay with that. I have limited time and barley understand the work!

Naciama:

hey!

SmokeyBrown:

I think, in general, if you write a radical equation that does not include negative roots, you should be able to avoid extraneous solutions?

SmokeyBrown:

Like the example above, √(x-2) = (x- 4) had the possibility of negative roots, in case x was less than 2? So, if you write a radical equation where that doesn't happen I'm guessing you should be good. Wanna give it a try, and we can see how it goes?

Naciama:

yeah we could do that. Im not top good at math so please bare with me

SmokeyBrown:

No problem, I'm no expert at this either, but we can try to work through it together

Naciama:

so how would you recomend we start

SmokeyBrown:

I guess one way we could start is by writing an equation that is true, like 9 = 9 Then we could think of a way to get a radical equation with variables out of that. So, let's say we have a variable x, which is our solution. I'll make it 7, for no particular reason. In that case, this is true: 9 = x + 2 As well as 9 = √(12x + 9) So maybe our equation could be √(12x + 9) = (x + 2) ?

Florisalreadytaken:

first of all, how familiar is the user with extraneous solies? for extraneous solie you could use \( \sqrt{x+2}=(0-3) \), where on the other hand for a not extraneos solie, just use a number for 0, just like \( \sqrt{x+2}=(8-3) \) work the examples, and check -- its more of a fixed thing, rather than one you can come up in the spot.

SmokeyBrown:

Well, we can square both sides and see if an extraneous solution results... So, from that original equation √(12x + 9) = (x + 2) both sides get squared to become 12x + 9 = x^2 + 4x + 4 Which we can simplify into 0 = x^2 - 8x -5 ..... It's already kind of looking like this will have more than one solution. Maybe I'm going about this wrong? I get the feeling that the solution has something to do with writing an equation that is a perfect square, since there will only be one solution to that. So perhaps we should start with that and move backwards?

SmokeyBrown:

@florisalreadytaken wrote:
first of all, how familiar is the user with extraneous solies? for extraneous solie you could use \( \sqrt{x+2}=(0-3) \), where on the other hand for a not extraneos solie, just use a number for 0, just like \( \sqrt{x+2}=(8-3) \) work the examples, and check -- its more of a fixed thing, rather than one you can come up in the spot.
tbh, no I don't really understand how extraneous solutions work, just the basic definitions. I appreciate you stepping in and taking a look!